# Albrechtstr. 30, 49076 Osnabrück, Germany

### v1

@@ Abstract
This empirical study aims at improving the effectiveness of the standard 2-layer MLP projection head featured in the SimCLR framework through the use of pretrained autoencoder embeddings. Given a contrastive learning task with a largely unlabeled image classification dataset, we first train a shallow autoencoder architecture and extract its compressed representations contained in the encoder’s embedding layer. After freezing the weights within this pretrained layer, we use it as a drop-in replacement for the input layer of SimCLR’s default projector. Additionally, we also apply further architectural changes to the projector by decreasing its width and changing its activation function. The different projection heads are then used to contrastively train and evaluate a feature extractor following the SimCLR protocol, while also examining the performance impact of -score normalized datasets. Our experiments indicate that using a pretrained autoencoder embedding in the projector can not only increase classification accuracy by up to 2.9 % or 1.7 % on average but can also significantly decrease the dimensionality of the projection space. Our results also suggest, that using the sigmoid and activation functions within the projector can outperform ReLU in terms of peak and average classification accuracy. When applying our presented projectors, then not applying -score normalization to datasets often increases peak performance. In contrast, the default projection head can benefit more from normalization. All experiments involving our pretrained projectors are conducted with frozen embeddings, since our test results indicate an advantage compared to using their non-frozen counterparts.
Keywords:
Nonlinear Projection Heads Multilayer Perceptrons Autoencoder Embeddings SimCLR Framework Contrastive Learning Representation Learning1 Introduction
@@
Contrastive learning is a self supervised machine learning paradigm that has proven itself to be useful for Big Data applications, which can capture nearly unlimited amounts of unlabeled data at high speeds and low costs. In such situations, classical supervised learning approaches become infeasible due to missing labels and because they can neither be generated in time nor in sufficient quality. For this reason, despite the groundbreaking success of supervision in the field of machine learning, the research community will continue to investigate new and existing methods for learning representation with a limited amount of labeled data. This paper combines the well understood fundamentals of unsupervised learning using autoencoders with more recent approaches of contrastive learning as offered by the SimCLR framework [2]. Both autoencoders and SimCLR are designed to be trained without any labeled data, although SimCLR’s evaluation protocol requires some amount of labeled samples. This is due to the fact that SimCLR is usually used for training image classification models, whereas autoencoders are often utilized for image generation and reconstruction tasks.
@@
In order to understand how the concepts behind SimCLR and plain autoencoders are combined in this paper, one first has to understand the abstract structure of SimCLR. At its most basic level, SimCLR is a composition of two artificial neural networks, the backbone followed by the projection head . Instead of focusing on the backbone, which is a deep convolutional neural network responsible for feature extraction, we direct our attention to the projection head, a shallow multilayer perceptron (MLP) consisting of two ReLU activated linear layers. The main role of this architectural building block is to project the features generated by the backbone into a lower dimensional projection space. This latent space and its embeddings can significantly influence the backbone’s ability to learn meaningful representations as the contrastive loss is applied to the outputs of the projector. This highlights the importance of the projection head within the SimCLR framework since backpropagation will always adapt the backbone’s weights based on the loss values generated by the compressed projections. Also, compressing certain inputs into more efficient representations can generally be seen as the central task of autoencoders. In fact, we find regular autoencoders to be a special type of MLP that can be trained to reconstruct inputs by optimizing a hidden embedding layer using some reconstruction loss that captures the difference between the encoder’s input and the decoder’s output. Based on these observations, we expect to increase the quality of contrastively learned representations by exchanging the randomly initialized input layer of SimCLR’s default projector with a pretrained autoencoder embedding.
@@
From a practical point of view, it is well known that deeper and wider backbones usually improve representation quality and downstream classification accuracy. In contrast, it is generally not well understood, why the backbone benefits from using a nonlinear projection head. Chen et al. [2] introduce the projector after empirically observing significant increases in model accuracy by more than 10 %. From this the authors conclude that the hidden layer before the projection head learns better representations than the layer after. The authors also conjecture that including a projection head increases the backbone’s ability to aggregate and retain information that is valuable for downstream tasks. However, the rational given for this phenomenon only explains why the backbone’s learned representations are of higher quality but not how the projector is able to generate them in the first place. This lack of reasoning becomes obvious when comparing the proposed end-to-end optimization of both the backbone and the projector to only training the backbone without any projection head. Despite some open questions about the projector’s precise mode of operation, its contribution to the success of SimCLR and related methods remains undeniable. We therefore conduct experiments as part of this paper to investigate the effectiveness of pretrained autoencoder embeddings within SimCLR’s default projector. The main contributions of our paper are summarized as follows:
@@
We demonstrate on five well-known image classification datasets that exchanging the input layer of SimCLR’s default projector with a pretrained autoencoder embedding reliably increases linear classification accuracy.
@@
We show that activating the projector with the sigmoid or functions can outperform ReLU in terms of peak and average accuracy.
@@
We observe that pretrained autoencoder embeddings can drastically reduce the width of the projector as well as the dimensionality of the projection space while increasing average accuracy.
4.
We find that freezing the weights within the pretrained embedding layers almost always increases classification accuracy.
5.
We note that -score normalization of datasets drastically impairs the training of autoencoder embeddings.
6.
To ensure reproducibility of our results, we publish the training and evaluation code together with CSV listings of all training runs on GitHub: https://github.com/andreas-schliebitz/simclr-ae
@@
The rest of the paper is organized as follows: Section 2 discusses related work. Next, we start to formally introduce our methods in Section 3, which contains an overview about the SimCLR framework, basic autoencoders, and our proposed architecture. In Section 4 we present and compare our results to SimCLR’s default projector and discuss them in Section 5. Finally, Section 6 concludes our paper by summarizing our work and giving interesting ideas for future research.
2 Related Work
@@
Literature review does not reveal any references to publications dealing explicitly with pretraining the nonlinear projection head used in the SimCLR framework. A few papers aim at explaining the role and the effects of the projector in more detail. These include the study conducted by Jing et al. [7] who are trying to understand dimensional collapse in contrastive self-supervised learning (SSL). The authors first show that dimensional collapse can also happen in contrastive learning despite the use of positive and negative samples. For SimCLR, dimensional collapse of the representation space usually occurs when the backbone is trained without a projection head (). Instead of improving the projection head itself, the authors propose an alternative to SimCLR called DirectCLR, a contrastive learning method that does not rely on a trainable projection head but uses a fixed low rank diagonal projector instead. DirectCLR does not try to prevent dimensional collapse by using a trainable projector but rather through direct optimization of the backbone. The proposed approach outperforms SimCLR when using a 1-layer linear projector by 1.6 % but fails to match the performance of SimCLR’s default 2-layer nonlinear projection head by 3.8 %.
@@
In contrast to Jing et al., the work of Gupta et al. [4] aims at understanding and improving the role of the projection head in self-supervised learning. The authors conjecture that the projection head implicitly learns to choose a subspace of features on which the contrastive loss is applied. This subspace selection is believed to address the disadvantages of the contrastive loss function which are mainly introduced by sub-optimal data augmentations. It is also assumed that implicit subspace selection can enable the projection head to generate embeddings which minimize contrastive loss, while allowing the backbone to learn meaningful representations. Based on these observations, the authors argue that data-dependent subspace selection should be part of the loss function used in self-supervised learning. This is why the authors reformulate SSL as a bilevel optimization problem where at each training step, the first optimization problem is to select the best subspace for the contrastive loss by optimizing the projection head while the second optimization performs gradient descent on the backbone. As a result, a contrastive loss function is proposed which treats the projector as part of the criterion and not only as an architectural component. The authors continue to empirically demonstrate that a trainable projector generally increases performance and can therefore be viewed as an improved optimization scheme for SSL. This result contradicts the general notion of DirectCLR [7], where backbone features are optimized directly using a non-trainable projector.
@@
Since contrastive learning is not only limited to SimCLR and its nonlinear projection head, Gui et al. [3] address the question why linear projectors affect the generalization properties of learned representations. The authors find the heuristics in the existing literature to be insufficient for answering this question, especially in the case of linear projection heads. Since discarding the projection head after training is a common approach in contrastive learning, the authors assume projectors to act as a buffer component that protects the representation space from distortions induced by loss minimization carried out on the outputs of the projector. For nonlinear projection heads, this assumption shows significant overlap with the conjecture made by Chen et al. [2] w. r. t. the superiority of the layer before the projector. This may be rooted in information loss triggered by contrastive loss: i. e. , statistical shrinkage and expansion, i. e. a reduction or increase in the effects of sampling variation induced by the contrastive loss on the projector, are highly correlated with downstream accuracy. Expansion decreases test error, while shrinkage reduces signals and increase test error.
@@
Xue et al. [9] theoretically investigate the benefits of a projection head for representation learning. The authors also observe that linear models progressively assign weights to features as they propagate through the layers of the backbone. Based on this observation, the authors state that features in deeper layers of the backbone are represented more unequally. Furthermore, these representations not only turn out to be more specialized towards the pretraining objective but are believed to be additionally distorted through the use of nonlinear activations within the layers of the backbone. The authors argue that this allows deeper layers to learn features which are entirely absent in the outputs of the projector. As a major result, it is stated that “projection heads provably improve the robustness and generalizability of representations, when data augmentation harms useful features of the pretraining data, or when features relevant to the downstream task are too weak or too strong in the pretraining data”. The authors show that these findings also apply to supervised contrastive learning and regular supervised learning where lower layers can also learn subclass-level features that are not represented in the final layer. Finally, it is noted that these findings “demonstrate how representations before the final representation layer can significantly reduce class/neural collapse”. In current literature, this effect is well observed for both nonlinear and linear projection heads [2], although DirectCLR [7] suggests linear projectors to be outdated.
@@
In summary, much of the current literature focuses on understanding the usefulness of projection heads for contrastive learning applications mainly from a theoretical point of view. We conclude from our literature review, that projection heads do increase the quality of contrastively learned representations in the backbone. Here, linear projection heads are significantly inferior to nonlinear projectors and that the advantage gained by linear projectors can also be experienced by directly optimizing the backbone. In contrast to linear projectors, there is no compelling evidence to suggest that nonlinear projection heads are redundant in contrastive learning applications. In fact, research indicates that nonlinear projectors continue to provide an easy and effective building block for achieving peak downstream performance while reducing the impact of model collapse.
3 Method
3.1 The SimCLR Framework
@@
SimCLR is a contrastive learning framework by Chen et al. [2]. It allows for learning representations from input images without the use of labeled data. The SimCLR framework builds upon extensively researched and carefully parameterized data augmentation pipelines used to create a positive contrastive image pair for each dataset sample . Here and are two transformations which are randomly sampled from the space of particularly effective transformations. These transformations include random cropping and flipping as well as the random application of color jitter, grayscaling and dynamic Gaussian blurring where the kernel size adapts to changes in image resolution.
@@
The composite and easy to implement architecture of SimCLR features an interchangeable feature extractor (backbone), followed by a ReLU activated 2-layer nonlinear projection head (projector). The backbone can be any neural network, although in practice, convolutional architectures like ResNet are widely used, especially for image classification tasks. Given an augmented pair of images , SimCLR first generates two separate intermediate representations by applying the feature extractor to both and . These representations are then mapped via the projector into a lower dimensional projection space:
@@
a variant of the InfoNCE loss [8], which does not require explicit sampling of negative examples for a given positive pair. Instead, given a batch of samples, all augmented samples (excluding the positive pair) are treated as negative examples. Research by Chen et al. [2] suggests, that this approach requires large batch sizes of up to 4096 to reach maximum efficacy. The NT-Xent loss shown in Equation 2 is calculated by taking the negative natural logarithm of the normalized exponential function (softmax), which is applied to the cosine similarities of the projections and scaled by a temperature parameter .
@@
3.2 Autoencoders
An autoencoder [6] is a type of multilayer perceptron consisting of an encoder and a decoder. The encoder can be formalized as a mapping from an input space into the embedding space , where both and are often euclidean spaces with . For an input vector the output of the encoder is called its embedding or latent vector . The decoder is usually the mirror image of the encoder. It maps the encoder’s outputs from the embedding space back into the space of decoded messages . The autoencoder can therefore be written as a function composition of followed by , mapping input via some to its reconstructed output :
@@ 3.2 Autoencoders
3.3 Pretraining Autoencoder Embeddings
The pretrained autoencoder embeddings used in this paper are generated by training a basic 4-layer autoencoder on five well-known image classification datasets (see Table 1) using gradient descent and MSE loss. For this part of our research, we use the base implementation of the autoencoder provided by Chadebec et al. [1] through the pythae Python library.
For each of the five datasets, we train three embeddings in a 70/10/20 train, validation and test split using varying latent dimensions of 128, 256 and 512. This equates to a total of 15 different pretrained autoencoder embeddings. Dataset samples are either used in their native resolution or scaled down to a maximum of px in order to reduce both the number of trainable parameters and GPU memory consumption. We do not apply -score normalization to the datasets as previous experiments with normalization resulted in validation loss becoming unresponsive.
All autoencoder embeddings are trained in parallel on three NVIDIA V100 GPUs with a fixed batch size of 100 and training times ranging from 25 to 50 minutes each, depending on dataset and embedding size. Each training run is configured analogous to Chadebec et al. [1, p. 25] to last for 100 epochs while using the Adam optimizer with an initial learning rate of and no weight decay. We parameterized the ReduceLROnPlateau learning rate scheduler to monitor the validation loss with a patience of 10 epochs and a reduction factor of 0.5. At the end of each training run, only the best performing autoencoder checkpoint is saved based on its validation loss.
3.4 Building Projectors from Autoencoder Embeddings

### v2

# An Empirical Study on Improving SimCLR’s Nonlinear Projection Head using Pretrained Autoencoder Embeddings

@@ Abstract
This paper focuses on improving the effectiveness of the standard 2-layer MLP projection head featured in the SimCLR framework through the use of pretrained autoencoder embeddings. Given a contrastive learning task with a largely unlabeled image classification dataset, we first train a shallow autoencoder architecture and extract its compressed representations contained in the encoder’s embedding layer. After freezing the weights within this pretrained layer, we use it as a drop-in replacement for the input layer of SimCLR’s default projector. Additionally, we also apply further architectural changes to the projector by decreasing its width and changing its activation function. The different projection heads are then used to contrastively train and evaluate a feature extractor following the SimCLR protocol. Our experiments indicate that using a pretrained autoencoder embedding in the projector can not only increase classification accuracy by up to 2.9 % or 1.7 % on average, but can also significantly decrease the dimensionality of the projection space. Our results also suggest, that using the sigmoid and tanh activation functions within the projector can outperform ReLU in terms of peak and average classification accuracy. All experiments involving our pretrained projectors are conducted with frozen embeddings, since our test results indicate an advantage compared to using their non-frozen counterparts.
I Introduction
@@
Contrastive learning is a self supervised machine learning paradigm that has proven itself to be useful for Big Data applications, which can capture nearly unlimited amounts of unlabeled data at high speed and low costs. Here, classical supervised learning approaches become infeasible due to missing labels, high training costs as well as often insufficient quality. For this reason, despite the groundbreaking success of supervision in the field of machine learning, the research community continues to investigate new and existing methods for learning representation with a limited amount of labeled data. This paper combines the well understood fundamentals of unsupervised learning using autoencoders with more recent approaches of contrastive learning as offered by the SimCLR framework [Chen2020]. Both autoencoders and SimCLR are designed to be trained without any labeled data, although SimCLR’s evaluation protocol requires some amount of labeled samples. This is due to the fact that SimCLR is usually used for training image classification models, whereas autoencoders are often utilized for image generation and reconstruction tasks.
@@
In order to understand how the concepts behind SimCLR and plain autoencoders are combined in this paper, one first has to understand the abstract structure of SimCLR. At its most basic level, SimCLR is a composition of two artificial neural networks, the backbone followed by the projection head . Instead of focusing on the backbone, which is a deep convolutional neural network responsible for feature extraction, we direct our attention to the projection head, a shallow multilayer perceptron (MLP) consisting of two ReLU activated linear layers. The main role of this MLP is to project the features generated by the backbone into a lower dimensional projection space. This latent space and its embeddings can significantly influence the backbone’s ability to learn meaningful representations as the contrastive loss is applied to the outputs of the projector. This highlights the importance of the projection head within the SimCLR framework since backpropagation will always adapt the backbone’s weights based on the loss values generated by the compressed projections. Also, compressing certain inputs into more efficient representations can generally be seen as the central task of autoencoders. In fact, we find regular autoencoders to be a special type of MLP that can be trained to reconstruct inputs by optimizing a hidden embedding layer using some reconstruction loss that captures the difference between the encoder’s input and the decoder’s output. Based on these observations, we expect to increase the quality of contrastively learned representations by exchanging the randomly initialized input layer of SimCLR’s default projector with a pretrained autoencoder embedding.
@@
From a practical point of view, it is well known that deeper and wider backbones usually improve representation quality and downstream classification accuracy [Szegedy2015, Zagoruyko16]. In contrast, it is generally not well understood, why the backbone benefits from using a nonlinear projection head. Chen et al. [Chen2020] introduce the projector after empirically observing significant increases in model accuracy by more than 10 %. From this, they conclude that the hidden layer before the projection head learns better representations than the layer after. The authors also conjecture that including a projection head increases the backbone’s ability to aggregate and retain information that is valuable for downstream tasks. However, the rationale given for this phenomenon only explains why the backbone’s learned representations are of higher quality but not how the projector is able to generate them in the first place. This lack of reasoning becomes obvious when comparing the proposed end-to-end optimization of both the backbone and the projector to only training the backbone without any projection head.
In general, despite some open questions about the projector’s precise mode of operation, its contribution to the success of SimCLR and related methods remains undeniable. We therefore conduct experiments as part of this paper in order to investigate the effectiveness of pretrained autoencoder embeddings within SimCLR’s default projector. The main contributions of our paper are summarized as follows:
@@
We analyze the impact of using a pretrained autoencoder within the SimCLR framework. In particular, we consider freezing the weights within the pretrained embedding layers and the influence of different activation functions (e. g., sigmoid or functions). We demonstrate on five well-known image classification datasets that exchanging the input layer of SimCLR’s default projector with a pretrained autoencoder embedding increases linear classification accuracy.
@@
We show, that pretrained autoencoder embeddings can considerably reduce the width of the projector and dimensionality of the projection space.
@@
We provide the code for our improved framework using pretrained autoencoder embeddings on GitHub (https://github.com/andreas-schliebitz/simclr-ae) to ensure the reproducibility of our results.
@@
The rest of the paper is organized as follows: Section II discusses related work. After that, Section III introduces the necessary background on SimCLR and Autoencoders. Next, we start to formally introduce our methods in Section IV, which contains an overview about the SimCLR framework, basic autoencoders, and our proposed architecture. In Section V we present and compare our results to SimCLR’s default projector and discuss them in Section VI. Finally, Section VII concludes our paper by summarizing our work and giving interesting ideas for future research.
II Related work
@@
Literature review does not reveal any references to publications dealing explicitly with pretraining the nonlinear projection head used in the SimCLR framework. A few papers aim at explaining the role and the effects of the projector in more detail. These include the study conducted by Jing et al. [Jing2021] who are trying to understand dimensional collapse in contrastive self-supervised learning (SSL). The authors first show that dimensional collapse can also happen in contrastive learning despite the use of positive and negative samples. For SimCLR, dimensional collapse of the representation space usually occurs when the backbone is trained without a projection head (). Instead of improving the projection head itself, the authors propose an alternative to SimCLR called DirectCLR, a contrastive learning method that does not rely on a trainable projection head but uses a fixed low rank diagonal projector instead. DirectCLR does not try to prevent dimensional collapse by using a trainable projector but rather through direct optimization of the backbone. The proposed approach outperforms SimCLR when using a 1-layer linear projector by 1.6 % but fails to match the performance of SimCLR’s default 2-layer nonlinear projection head by 3.8 %.
@@
In contrast to Jing et al. [Jing2021], the work of Gupta et al. [Gupta2022] aims at understanding and improving the role of the projection head in self-supervised learning. The authors conjecture that the projection head implicitly learns to choose a subspace of features on which the contrastive loss is applied. This subspace selection is believed to address the disadvantages of the contrastive loss function, mainly introduced by sub-optimal data augmentations. It is also assumed that implicit subspace selection can enable the projection head to generate embeddings minimizing contrastive loss, while allowing the backbone to learn meaningful representations. Based on these observations, the authors argue that data-dependent subspace selection should be part of the loss function used in self-supervised learning. This is why the authors reformulate SSL as a bilevel optimization problem where at each training step, the first optimization problem is to select the best subspace for the contrastive loss by optimizing the projection head while the second optimization performs gradient descent on the backbone. As a result, a contrastive loss function is proposed which treats the projector as part of the criterion and not only as an architectural component. The authors continue to empirically demonstrate that a trainable projector generally increases performance and can therefore be viewed as an improved optimization scheme for SSL. This result contradicts the general notion of DirectCLR [Jing2021], where backbone features are optimized directly using a non-trainable projector.
@@
Since contrastive learning is not only limited to SimCLR and its nonlinear projection head, Gui et al. [Gui2023] address the question why linear projectors affect the generalization properties of learned representations. The authors find the heuristics in the existing literature to be insufficient for answering this question, especially in the case of linear projection heads. Since discarding the projection head after training is a common approach in contrastive learning, the authors assume projectors to act as a buffer component that protects the representation space from distortions induced by loss minimization carried out on the outputs of the projector. For nonlinear projection heads, this assumption shows significant overlap with the conjecture made by Chen et al. [Chen2020] w. r. t. the superiority of the layer before the projector. This may be rooted in information loss triggered by contrastive loss: i. e. , statistical shrinkage and expansion, i. e. a reduction or increase in the effects of sampling variation induced by the contrastive loss on the projector, are highly correlated with downstream accuracy. Expansion decreases test error, while shrinkage reduces signals and increase test error.
@@
Xue et al. [Xue2024] theoretically investigate the benefits of a projection head for representation learning. The authors also observe that linear models progressively assign weights to features as they propagate through the layers of the backbone. Based on this observation, the authors state that features in deeper layers of the backbone are represented more unequally. Furthermore, these representations not only turn out to be more specialized towards the pretraining objective but are believed to be additionally distorted through the use of nonlinear activations within the layers of the backbone. The authors argue that this allows deeper layers to learn features which are entirely absent in the outputs of the projector. As a major result, it is stated that “projection heads provably improve the robustness and generalizability of representations, when data augmentation harms useful features of the pretraining data, or when features relevant to the downstream task are too weak or too strong in the pretraining data”. The authors show that these findings also apply to supervised contrastive learning and regular supervised learning where lower layers can also learn subclass-level features that are not represented in the final layer. Finally, it is noted that these findings “demonstrate how representations before the final representation layer can significantly reduce class/neural collapse”. In current literature, this effect is well observed for both nonlinear and linear projection heads [Chen2020], although DirectCLR [Jing2021] suggests linear projectors to be outdated.
@@
In summary, much of the current literature focuses on understanding the usefulness of projection heads for contrastive learning applications mainly from a theoretical point of view. We conclude from our literature review, that projection heads do increase the quality of contrastively learned representations in the backbone. Here, linear projection heads are significantly inferior to nonlinear projectors and that the advantage gained by linear projectors can also be experienced by directly optimizing the backbone. In contrast to linear projectors, there is no compelling evidence to suggest that nonlinear projection heads are ineffective in contrastive learning applications. In fact, research indicates that nonlinear projectors continue to provide an easy and effective building block for achieving peak downstream performance while reducing the impact of model collapse.
III Background
III-A The SimCLR framework
@@
SimCLR is a contrastive learning framework by Chen et al. [Chen2020]. It allows for learning representations from input images without the use of labeled data. The SimCLR framework builds upon extensively researched and carefully parameterized data augmentation pipelines used to create a positive contrastive image pair for each dataset sample . Here and are two transformations which are randomly sampled from the space of particularly effective transformations. These transformations include random cropping and flipping as well as the random application of color jitter, grayscaling and dynamic Gaussian blurring where the kernel size adapts to changes in image resolution.
@@
The composite and easy to implement architecture of SimCLR features an interchangeable feature extractor (backbone), followed by a ReLU [Nair2010] activated 2-layer nonlinear projection head (projector). The backbone can be any neural network, although in practice, convolutional architectures like ResNet [He2016] are widely used, especially for image classification tasks. Given an augmented pair of images , SimCLR first generates two separate intermediate representations by applying the feature extractor to both and . These representations are then mapped via the projector into a lower dimensional projection space:
@@
a variant of the InfoNCE loss [Oord2018], which does not require explicit sampling of negative examples for a given positive pair. Instead, given a batch of samples, all augmented samples (excluding the positive pair) are treated as negative examples. Research by Chen et al. [Chen2020] suggests, that this approach requires large batch sizes of up to 4096 to reach maximum efficacy. The NT-Xent loss shown in Equation 2 is calculated by taking the negative natural logarithm of the normalized exponential function (softmax), which is applied to the cosine similarities of the projections and scaled by a temperature parameter .
@@
III-B Autoencoders
An autoencoder [Hinton2006] is a type of multilayer perceptron consisting of an encoder and a decoder. The encoder can be formalized as a mapping from an input space into the embedding space , where both and are often euclidean spaces with . For an input vector the output of the encoder is called its embedding or latent vector . The decoder is usually the mirror image of the encoder. It maps the encoder’s outputs from the embedding space back into the space of decoded messages . The autoencoder can therefore be written as a function composition of followed by , mapping input via some to its reconstructed output :
@@ III-B Autoencoders
IV Method
IV-A Pretraining autoencoder embeddings
The pretrained autoencoder embeddings used in this paper are generated by training a basic 4-layer autoencoder on five well-known image classification datasets (see Table I) using gradient descent and MSE loss. For this part of our research, we use the base implementation of the autoencoder provided by Chadebec et al. [Chadebec2022] through the pythae Python library.
For each of the five datasets, we train three embeddings in a 70/10/20 train, validation and test split using varying latent dimensions of 128, 256 and 512. This equates to a total of 15 different pretrained autoencoder embeddings. Dataset samples are either used in their native resolution or scaled down to a maximum of px in order to reduce both the number of trainable parameters and GPU memory consumption.
All autoencoder embeddings are trained in parallel on three NVIDIA V100 GPUs with a fixed batch size of 100 and training times ranging from 25 to 50 minutes each, depending on dataset and embedding size. Each training run is configured analogous to Chadebec et al. [Chadebec2022, p. 25] to last for 100 epochs while using the Adam optimizer [Kingma14] with an initial learning rate of and no weight decay. We parameterized the ReduceLROnPlateau learning rate scheduler to monitor the validation loss with a patience of 10 epochs and a reduction factor of 0.5. At the end of each training run, only the best performing autoencoder checkpoint is saved based on its validation loss.
IV-B Building projectors from autoencoder embeddings